A​ one-year discount bond with a face value of​ $1,000 that is currently selling for​ $900 has an interest rate of A. ​5.26%. B. ​10%. C. ​11.1%. D. ​100%.
Answer #1

PV = FV / (1 + r)n

r = [FV / PV]1/n - 1

= [$1,000 / $900] - 1 = 1.111 - 1 = 0.111, or 11.1%

Hence, Option "C" is correct.